(1+y2)dx=(tan-1 y-x)dy

1
Expert's answer
2021-06-07T13:27:08-0400

"\text{Given, the differential equation,}\\
(1+y^2)dx=(tan^{-1} y-x)dy\\
\frac{dx}{dy}=\frac{tan^{-1}y}{1+y^2}-\frac{x}{1+y^2}\\
\frac{dx}{dy}+\frac{x}{1+y^2}
=\frac{tan^{-1}y}{1+y^2}\\
\text{This is a linear differential equation in term of x.}\\
\text{
Integrating factor, }\\
IF=e^{ \int \frac{dy}{1+y^2}} \\
\text{Then, solution is given by} \\
xe^{tan^{-1}y}= \int \frac{tan^{-1}y}{1+y^2}e^{tan^{-1}y}dy\\
Put \\
tan^{-1}y=t\implies \frac{dy}{1+y^2}=dt\\
Therefore,\\
\int \frac{tan^{-1}y}{1+y^2}e^{tan^{-1}y}dy=-\int te^{t}dt\\
=(1-t)e^t\\
=(1-tan^{-1}y)e^{tan^{-1}y}+C\\
Then, the\space solution\\
xe^{tan^{-1}y}=(1-tan^{-1}y)e^{tan^{-1}y}+C\\"